what is the perimeter,in meters.of a rectangular play groung 24 meters wide that has a same area as a rectangular play groung 64 meters long and 48 meters wide.
Answer by jbevis123080(1)

304 meters.
First find the area of the playground dimensions you know (64m by 48m). Area=length X width, so Area of playground #2 = 64X48=3072 meters squared. Then go back to your other playground and divide 3072 by 24 = 128m. Perimeter of a rectangle = 2(length) + 2(width), so the perimeter = 2(128)+2(24)= 256+48 = 304 meters